# Two-Stage Nuisance Function Estimation for Causal Mediation Analysis

### v2

@@ Appendix A Proofs
    $\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]$
    $=\mathbb{E}\left[\left\{\frac{p(A=1\mid M,X)p(M\mid A=0,X)}{p(M,A=1\mid X)}-\frac{p(A=0\mid M,X)}{p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
    $=\mathbb{E}\left[\left\{\frac{p(M\mid A=0,X)}{p(M\mid X)}-\frac{p(A=0\mid M,X)}{p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
    $=\mathbb{E}\left[\left\{\frac{p(M,A=0\mid X)}{p(M\mid X)p(A=0\mid X)}-\frac{p(M,A=0\mid X)}{p(M\mid X)p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
@@
$\mathbb{E}[\hat{\psi}^{mr}]-\psi_{0}$
$=\mathbb{E}[\hat{\mu}_{2}(X)+(1-A)\hat{\pi}_{1}(X)\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{2}(X)\}+A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}]-\psi_{0}$
$=\mathbb{E}\left[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}~{}~{}+\mathbb{E}\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\left\{\mathbb{E}\left[1-A\mid X\right]\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}~{}~{}+\mathbb{E}\left[\left\{\mathbb{E}\left[A\mid M,X\right]\hat{\pi}_{2}(M,X)-\mathbb{E}\left[1-A\mid M,X\right]\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\left\{p\left(A=0\mid X\right)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}~{}~{}+\mathbb{E}\left[\left\{p\left(A=1\mid M,X\right)\hat{\pi}_{2}(M,X)-p\left(A=0\mid M,X\right)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
@@
$\mathbb{E}[\hat{\theta}^{mr}]-\theta_{0}$
$=\mathbb{E}\left[\hat{\mu}_{2}(X)+(1-A)\hat{\pi}_{1}(X)\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{2}(X)\}+A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}\right]-\mathbb{E}[\mu_{2}(X)]$
$=\mathbb{E}\left[\hat{\mu}_{2}(X)-\mu_{2}(X)+(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\mathbb{E}\left[(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}\bigg{|}X\right]-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\mathbb{E}\left[A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}\bigg{|}M,X\right]-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\mathbb{E}[(1-A)\mu_{1}(M,X)\mid X]\hat{\pi}_{1}(X)-\frac{\hat{\pi}_{1}(X)\hat{\mu}_{2}(X)}{\pi_{1}(X)}-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\Big{[}\hat{\pi}_{2}(M,X)\big{(}\mathbb{E}[AY\mid M,X]-\mathbb{E}[A\mid M,X]\hat{\mu}_{1}(M,X)\big{)}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\Big{]}$
$=\mathbb{E}\left[\mathbb{E}[(1-A)\mid X]\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\hat{\pi}_{1}(X)-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\hat{\pi}_{2}(M,X)\mathbb{E}[A\mid M,X]\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
@@
Assume the size of each fold of the data is $m$, that is, $n=mL$. Let $(\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})$ be the estimators of $(\pi_{1},\pi_{2},\mu_{1},\mu_{2})$ obtained on subsamples $I_{\ell}^{c}$. Then $\hat{\psi}^{mr}_{\ell}$ can be written as
$\hat{\psi}^{mr}_{\ell}=\mathbb{E}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right].$
@@
$\sqrt{n}(\hat{\psi}^{mr}-\psi_{0})$ $=\sqrt{n}\Big{\{}\frac{1}{L}\sum_{\ell=1}^{L}\mathbb{E}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big{\}}$
@@
$\quad\quad\quad\quad\quad\quad\quad+\sqrt{m}\Big{\{}\mathbb{E}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big{\}}\Bigg{\}},$
@@
$\mathbb{G}_{n}[f]=\sqrt{n}\{\mathbb{E}_{n}[f]-\mathbb{E}[f]\},$
@@
$\sqrt{n}(\hat{\psi}^{mr}-\psi_{0})$ $=\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\left\{\mathbb{G}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]\right\}$ $(T_{1})$
@@
$\quad+\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\sqrt{m}\Big{\{}\mathbb{E}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big{\}}.$ $(T_{3})$
We will show that $(T_{2})$ provides us with the term $\sqrt{n}\mathbb{E}_{n}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]$, and under the assumptions of the theorem, $(T_{1})$ and $(T_{3})$ are $o_{p}(1)$.
Analysis of $(T_{1})$ Define
@@
$=m\cdot var\left(\mathbb{E}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\mathbb{E}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]\mid I_{\ell}^{c}\right)$
@@
$=\mathbb{E}\Big{[}\Big{\{}\hat{\mu}_{2,\ell}(X)\left(1-(1-A)\hat{\pi}_{1,\ell}(X)\right)+\hat{\mu}_{1,\ell}(M,X)\left((1-A)\hat{\pi}_{1,\ell}(X)-A\hat{\pi}_{2,\ell}(M,X)\right)+AY\hat{\pi}_{2,\ell}(M,X)$
@@
$\mathbb{E}\left[A_{m}^{\ell}\mid I_{\ell}^{c}\right]$
$=\mathbb{E}\Big{[}\hat{\mu}_{2,\ell}(X)+(1-A)\hat{\pi}_{1,\ell}(X)\{\hat{\mu}_{1,\ell}(M,X)-\hat{\mu}_{2,\ell}(X)\}+A\hat{\pi}_{2,\ell}(M,X)\{Y-\hat{\mu}_{1,\ell}(M,X)\}$
@@
$=\mathbb{E}\Big{[}\hat{\mu}_{2,\ell}(X)\left(1-(1-A)\hat{\pi}_{1,\ell}(X)\right)+\hat{\mu}_{1,\ell}(M,X)\left((1-A)\hat{\pi}_{1,\ell}(X)-A\hat{\pi}_{2,\ell}(M,X)\right)+AY\hat{\pi}_{2,\ell}(M,X)$
@@
$\quad\quad+\Big{\|}\hat{\mu}_{1,\ell}^{2}(M,X)\Big{\|}_{2}^{1/2}\Big{\|}\big{(}\hat{\pi}_{1,\ell}(X)-\pi_{1}(X)\big{)}^{2}\Big{\|}_{2}^{1/2}$
$\quad\quad+\Big{\|}\hat{\mu}_{1,\ell}^{2}(M,X)\Big{\|}_{2}^{1/2}\Big{\|}\big{(}\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X)\big{)}^{2}\Big{\|}_{2}^{1/2}$
$\quad\quad+\Big{\|}\mu_{1}^{2}(M,X)\Big{\|}_{2}^{1/2}\Big{\|}(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))^{2}\Big{\|}_{2}^{1/2}.$
@@
$\mathbb{E}\left[(A_{m}^{\ell})^{2}\mid I_{\ell}^{c}\right]=var\left(A_{m}^{\ell}\mid I_{\ell}^{c}\right)+\Big{(}\mathbb{E}\left[A_{m}^{\ell}\mid I_{\ell}^{c}\right]\Big{)}^{2}\stackrel{{\scriptstyle p.}}{{\rightarrow}}0,\quad\text{ as }m\rightarrow\infty.$
@@
$\mathbb{E}\left[P(|A_{m}^{\ell}|>\delta\mid I_{\ell}^{c})\right]=P(|A_{m}^{\ell}|>\delta)\stackrel{{\scriptstyle p.}}{{\rightarrow}}0,\quad\text{ as }m\rightarrow\infty.$
@@
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\left\{\mathbb{E}_{m}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]-\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]\right\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\frac{1}{m}\sum_{\ell=1}^{L}\sum_{i=1}^{m}\left\{\phi(O_{i};\pi_{1},\pi_{2},\mu_{1},\mu_{2})-\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]\right\}$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left\{\phi(O_{i};\pi_{1},\pi_{2},\mu_{1},\mu_{2})-\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]\right\}.$
By central limit theorem, the last expression converges to the distribution $\mathcal{N}(0,var(\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2}))$ if $\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]<\infty$. Note that $\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]$ can be upper bounded as
$\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]$
$=\mathbb{E}\Big{[}\Big{\{}\mu_{2}(X)+(1-A)\pi_{1}(X)\{\mu_{1}(M,X)-\mu_{2}(X)\}+A\pi_{2}(M,X)\{Y-\mu_{1}(M,X)\}\Big{\}}^{2}\Big{]}$
$=\mathbb{E}\Big{[}\Big{\{}\mu_{2}(X)\left(1-(1-A)\pi_{1}(X)\right)+\mu_{1}(M,X)\left((1-A)\pi_{1}(X)-A\pi_{2}(M,X)\right)+AY\pi_{2}(M,X)\Big{\}}^{2}\Big{]}$
@@
Therefore, by Assumption 3 (c), we have $\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]<\infty$.

### v3

@@ Appendix A Proofs
    $E\left[\left\{A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]$
    $=E\left[\left\{\frac{p(A=1\mid M,X)p(M\mid A=0,X)}{p(M,A=1\mid X)}-\frac{p(A=0\mid M,X)}{p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
    $=E\left[\left\{\frac{p(M\mid A=0,X)}{p(M\mid X)}-\frac{p(A=0\mid M,X)}{p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
    $=E\left[\left\{\frac{p(M,A=0\mid X)}{p(M\mid X)p(A=0\mid X)}-\frac{p(M,A=0\mid X)}{p(M\mid X)p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
@@
$E[\hat{\psi}^{\text{2S}}]-\psi_{0}$
$=E[\hat{\mu}_{2}(X)+(1-A)\hat{\pi}_{1}(X)\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{2}(X)\}+A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}]-\psi_{0}$
$=E\left[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~~~+E\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=E\left[\left\{E\left[1-A\mid X\right]\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~~~+E\left[\left\{E\left[A\mid M,X\right]\hat{\pi}_{2}(M,X)-E\left[1-A\mid M,X\right]\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=E\left[\left\{p\left(A=0\mid X\right)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~~~+E\left[\left\{p\left(A=1\mid M,X\right)\hat{\pi}_{2}(M,X)-p\left(A=0\mid M,X\right)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
@@
$E[\hat{\psi}^{\text{2S}}]-\psi_{0}$
$=E\left[\hat{\mu}_{2}(X)+(1-A)\hat{\pi}_{1}(X)\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{2}(X)\}+A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}\right]-E[\mu_{2}(X)]$
$=E\left[\hat{\mu}_{2}(X)-\mu_{2}(X)+(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\left[A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=E\left[(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\left[A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=E\left[E\left[(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}\bigg|X\right]-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\left[E\left[A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}\bigg|M,X\right]-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=E\left[E[(1-A)\mu_{1}(M,X)\mid X]\hat{\pi}_{1}(X)-\frac{\hat{\pi}_{1}(X)\hat{\mu}_{2}(X)}{\pi_{1}(X)}-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\Big[\hat{\pi}_{2}(M,X)\big(E[AY\mid M,X]-E[A\mid M,X]\hat{\mu}_{1}(M,X)\big)-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\Big]$
$=E\left[E[(1-A)\mid X]\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\hat{\pi}_{1}(X)-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\left[\hat{\pi}_{2}(M,X)E[A\mid M,X]\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=E\left[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
@@
Assume the size of each fold of the data is $m$, that is, $n=mL$. Let $(\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})$ be the estimators of $(\pi_{1},\pi_{2},\mu_{1},\mu_{2})$ obtained on subsamples $I_{\ell}^{c}$. Then $\hat{\psi}^{\text{2S}}_{\ell}$ can be written as
$\hat{\psi}^{\text{2S}}_{\ell}=E_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right].$
@@
$\sqrt{n}(\hat{\psi}^{\text{2S}}-\psi_{0})$ $=\sqrt{n}\Big\{\frac{1}{L}\sum_{\ell=1}^{L}E_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big\}$
@@
$\quad\quad\quad\quad\quad\quad\quad+\sqrt{m}\Big\{E\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big\}\Bigg\},$
@@
$\mathbb{G}_{n}[f]=\sqrt{n}\{E_{n}[f]-E[f]\},$
@@
$\sqrt{n}(\hat{\psi}^{\text{2S}}-\psi_{0})$ $=\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\left\{\mathbb{G}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]\right\}$ $(T_{1})$
@@
$\quad+\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\sqrt{m}\Big\{E\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big\}.$ $(T_{3})$
We will show that $(T_{2})$ provides us with the term $\sqrt{n}E_{n}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]$, and under the assumptions of the theorem, $(T_{1})$ and $(T_{3})$ are $o_{p}(1)$.
Analysis of $T_{1}$
Define
@@
$=m\cdot var\left(E_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-E_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]\mid I_{\ell}^{c}\right)$
@@
$=E\Big[\Big\{\hat{\mu}_{2,\ell}(X)\left(1-(1-A)\hat{\pi}_{1,\ell}(X)\right)+\hat{\mu}_{1,\ell}(M,X)\left((1-A)\hat{\pi}_{1,\ell}(X)-A\hat{\pi}_{2,\ell}(M,X)\right)+AY\hat{\pi}_{2,\ell}(M,X)$
@@
$E\left[A_{m}^{\ell}\mid I_{\ell}^{c}\right]$
$=E\Big[\hat{\mu}_{2,\ell}(X)+(1-A)\hat{\pi}_{1,\ell}(X)\{\hat{\mu}_{1,\ell}(M,X)-\hat{\mu}_{2,\ell}(X)\}+A\hat{\pi}_{2,\ell}(M,X)\{Y-\hat{\mu}_{1,\ell}(M,X)\}$
@@
$=E\Big[\hat{\mu}_{2,\ell}(X)\left(1-(1-A)\hat{\pi}_{1,\ell}(X)\right)+\hat{\mu}_{1,\ell}(M,X)\left((1-A)\hat{\pi}_{1,\ell}(X)-A\hat{\pi}_{2,\ell}(M,X)\right)+AY\hat{\pi}_{2,\ell}(M,X)$
@@
$\leq\Big\|\hat{\mu}_{2,\ell}^{2}(X)\Big\|_{2}^{1/2}\Big\|(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X))^{2}\Big\|_{2}^{1/2}+\Big\|\hat{\mu}_{1,\ell}^{2}(M,X)\Big\|_{2}^{1/2}\Big\|\big(\hat{\pi}_{1,\ell}(X)-\pi_{1}(X)\big)^{2}\Big\|_{2}^{1/2}$
$\quad\quad+\Big\|\hat{\mu}_{1,\ell}^{2}(M,X)\Big\|_{2}^{1/2}\Big\|\big(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X)\big)^{2}\Big\|_{2}^{1/2}+\Big\|\mu_{1}^{2}(M,X)\Big\|_{2}^{1/2}\Big\|(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))^{2}\Big\|_{2}^{1/2}.$
@@
$E\left[(A_{m}^{\ell})^{2}\mid I_{\ell}^{c}\right]=var\left(A_{m}^{\ell}\mid I_{\ell}^{c}\right)+\Big(E\left[A_{m}^{\ell}\mid I_{\ell}^{c}\right]\Big)^{2}\stackrel{{\scriptstyle p.}}{{\rightarrow}}0,\quad\text{ as }m\rightarrow\infty.$
@@
$E\left[P(|A_{m}^{\ell}|>\delta\mid I_{\ell}^{c})\right]=P(|A_{m}^{\ell}|>\delta)\stackrel{{\scriptstyle p.}}{{\rightarrow}}0,\quad\text{ as }m\rightarrow\infty.$
@@
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\left\{E_{m}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]-E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]\right\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\frac{1}{m}\sum_{\ell=1}^{L}\sum_{i=1}^{m}\left\{\phi(O_{i};\pi_{1},\pi_{2},\mu_{1},\mu_{2})-E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]\right\}$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left\{\phi(O_{i};\pi_{1},\pi_{2},\mu_{1},\mu_{2})-E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]\right\}.$
By central limit theorem, the last expression converges to the distribution $\mathcal{N}(0,var(\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2}))$ if $E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]<\infty$. Note that $E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]$ can be upper bounded as
$E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]$
$=E\Big[\Big\{\mu_{2}(X)+(1-A)\pi_{1}(X)\{\mu_{1}(M,X)-\mu_{2}(X)\}+A\pi_{2}(M,X)\{Y-\mu_{1}(M,X)\}\Big\}^{2}\Big]$
$=E\Big[\Big\{\mu_{2}(X)\left(1-(1-A)\pi_{1}(X)\right)+\mu_{1}(M,X)\left((1-A)\pi_{1}(X)-A\pi_{2}(M,X)\right)+AY\pi_{2}(M,X)\Big\}^{2}\Big]$
@@
Therefore, by Assumption 3 (c), we have $E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]<\infty$.